1)  Let us assume that from a population with mean  and standard deviation  a sample random variable of  is selected.

 

a)   What is the probability ?

 

b)  What is the probability ?

 

c)   What is the probability ?

 


1
Expert's answer
2021-08-16T15:30:52-0400

Given "\mu=100, \sigma=15, n=900"

"X\sim N(\mu, \sigma^2/n)"

a.

"P(X<101.1)=P(Z<\dfrac{101.1-100}{15/\sqrt{900}})"


"=P(Z<2.2)\approx0.9861"

b.

"P(X>101.5)=1-P(Z\leq\dfrac{101.5-100}{15/\sqrt{900}})"


"=1-P(Z\leq3)\approx0.0013"

c.

"P(99.3<X<100.5)=P(Z<\dfrac{100.5-100}{15/\sqrt{900}})"

"-P(Z\leq\dfrac{99.3-100}{15/\sqrt{900}})"

"=P(Z<1)-P(Z\leq-1.4)"


"\approx0.84134-0.08076\approx0.7606"